# Accelerating Ray Tracing-Based Wireless Channels Generation for Real-Time Network Digital Twins

### v1

@@ V-B2 Cut-out-based simplification methods
  • Coverage Map Method: also requires an RT simulation and removes regions based on a predefined path loss threshold (in $\mathrm{d}\mathrm{B}$). The scenario is pruned by discarding areas where the path loss exceeds the threshold, optimizing the input for subsequent RT simulations.
@@ VI Evaluation
All experiments were performed using Sionna RT v0.19.20.19.20.19.20.19.2 [47] in different computers. For evaluating solutions to P1 (WCG methods), the server was equipped with an Intel® Core i7-10700F CPU @ 2.90GHz, 128 GiB RAM. For the SOH methods (P2), the server was equipped with Intel® Xeon Gold 5419Y, 32 GiB RAM. In terms of simulation parameters, we considered a carrier frequency $f=2.14$ $\mathrm{G}\mathrm{H}\mathrm{z}$ and a narrowband spatial 2D channel model [21, 22, 7] defined by
$\mathbf{H}_{n}=\sum\limits_{r=1}^{K_{n}}\kappa_{r}\mathbf{a}_{\text{R}}(\theta_{r}^{\textrm{A}})\mathbf{a}_{\text{T}}^{*}(\theta_{r}^{\textrm{D}}),$ (5)
where $\mathbf{H}_{n}$ is the channel matrix for scene $n$, $\mathbf{a}_{\mathbf{T}}$, $\mathbf{a}_{\mathbf{R}}$ are the steering vectors at the transmitter (Tx) and receiver (Rx); and $\kappa_{r}$ is the $r$-th channel complex-valued gain for the $r$-th ray given by
$\kappa_{r}=\alpha_{r}e^{-j\Omega_{r}},$ (6)
where $\alpha_{r}$ and $\Omega_{r}$ are listed in Table I. The phase $\Omega_{r}$ is the most difficult parameter to predict among the ones listed in Table I. It can be written as
$\Omega_{r}=\psi_{r}-2\pi f\tau_{r},$ (7)
where $\psi_{r}$ is the phase parcel that depends on the ray interactions and the electromagnetic properties (e. g., permittivity, conductivity, permeability, and scattering coefficient) of the respective materials [7], and $\tau_{r}$ is the $r$-th path delay as listed in Table I.
Furthermore, we adopted a uniform linear array (ULA) MIMO system configuration with eight transmit antennas and four receive antennas, all featuring isotropic radiation patterns and cross-polarization. Additionally, all scenarios assume single mobility, with only the Rx moving along a linear track. Table III summarizes these and other simulation parameters used to validate the proposed methods in this example.
TABLE III: Simulation parameters used to evaluate the WCG module in each scenario
Simulation parameter Value
@@
Number of antennas per Tx $8$
Number of antennas per Rx $4$
@@
In order to evaluate the solutions to Problems P1 and P2, we consider each solution independently, in spite of the fact that in practice they can be combined by the WCG and SOH modules.
VI-A RT Augmentation (P1)
@@
$\hat{\mathbf{H}}_{p}=\mathbf{H}_{1}+\dfrac{p-1}{U}(\mathbf{H}_{2}-\mathbf{H}_{1}).$ (9)
To compare the performance of the ARTS method with this baseline in solving P1, we used three different urban macro scenarios: Munich, Etoile, and St. Canyon [47]. For all of them, the Tx height is greater than 25 $\mathrm{m}$. In terms of mobility, for all Rx, we consider a constant velocity of 3 $\mathrm{k}\mathrm{m}\mathrm{/}\mathrm{h}$. The simulations consider only reflection and diffraction, and two different upsampling factors: $U=2$ and $U=10$. The value $U=2$ is the smallest upsampling factor, while $U=10$ is much more challenging. Given $U$, the sampling interval $T_{o}$, and the receiver speed, one can estimate the number of meters the receiver moved between neighboring scenes. The conventional RT was executed for a total of $1000$ scenes, which are then processed and split into “known” channels and the ones that play the role of ground truth to assess augmentation. Table IV summarizes the experiments.
@@
# experiment $U$ Known scenes Known scenes density ($\mathrm{m}$) $T_{\textrm{o}}$ $T_{\textrm{s}}$
$1$ $2$ $500$ $1$ scene every $0.2$ $\mathrm{m}$ $0.24$ $0.12$
$2$ $10$ $100$ $1$ scene every $2$ $\mathrm{m}$ $2.4$ $0.24$
TABLE V: Comparison of results with ARTS and the baseline for Experiments 1 and 2 (upsampling $U=2$ and $U=10$, respectively)
@@
Experiment 1 Matrix linear interp. $-1.40$ $\mathbf{5.56}$ $-8.53$ $\mathbf{3.93}$ $-14.15$ $4.62$
Proposed ARTS method $\mathbf{-56.44}$ $6.28$ $\mathbf{-67.42}$ $4.69$ $\mathbf{-75.93}$ $\mathbf{4.55}$
Experiment 2 Matrix linear interp. $-1.33$ $\mathbf{6.11}$ $-2.46$ $5.95$ $-5.16$ $11.31$
@@
In Experiment 1, $500$ channels are known, provided by the ray tracer, and augmentation is used to estimate the remaining $500$ channels. The results are presented in Fig. 10, where the cumulative distribution functions show that ARTS outperforms interpolation in the matrix domain.
Figure 10: NMSE CDF considering our proposed method for Problem P1 and matrix interpolation at Munich, Etoile, and St. Canyon scenario.
As mentioned, Experiment 2 is more challenging given that $U=10$. Specifically, it required generating 900 scenes from information provided by 100 known scenes. The results in Fig. 11 illustrate the performance under this configuration.
Figure 11: NMSE CDF considering our proposed method and matrix interpolation at Munich, Etoile, and St. Canyon scenario with a fewer number of known samples.
The results in Fig. 11 indicate the expected decline in performance with respect to Fig. 10. This degradation can be attributed to the lack of spatial consistency between the RT samples. Specifically, the probability of obtaining two equivalent rays separated by 2 $\mathrm{m}$ is significantly lower compared to when the samples are only $0.2$ $\mathrm{m}$ apart. Despite this decline in NMSE, the proposed ARTS method outperforms the matrix interpolation method in all three scenarios, which is summarized in Table V. ARTS achieves a reduction of more than 17 $\mathrm{d}\mathrm{B}$ in average NMSE, while decreasing the maximum distortion in half of the cases.
@@
In order to validate solutions to P2 (the SOH methods) regarding simplification problems, we defined two experiments considering one more metric: the RT simulation time in comparison with that obtained in the original scenario. In the first experiment, we separately assessed all 3D simplification methods. In the second experiment, we evaluated the performance taking into account all possible combinations of mesh and cut-based simplification approaches. For both experiments, we considered a fourth scenario called Modern City, which is depicted in Fig. 12. Modern City has a larger number of meshes when compared to the three scenarios (Munich, Etoile, and St. Canyon) used to validate the WCG module. More specifically, Modern City has 5 922 924 faces; while Munich, Etoile, and St. Canyon have 38 938, 13 058, and 74 faces, respectively.
Figure 12: Modern City scenario used for evaluating the methods to solving P2.
The two experiments described in this subsection used the same ones mentioned in Table III. One exception is that, while the P1 solutions used only reflection and diffraction, here we also consider diffuse scattering. This increases the number of detected paths and, consequently, the complexity of RT, leading to more evident results for both NMSE and RT simulation times. Furthermore, the scope of the two new experiments comprises six simplification methods, all defined in Fig. 7. Specific parameters for each method are summarized in Table VI.
@@
To compute the NMSE, for each simplification method, we ran 150 scenes, varying the Rx position considering a pre-defined position path, as shown in Fig. 13. After that, we took the CDF, the mean, and the standard deviation (SD) between the values obtained. Thus, for this first experiment regarding P2, we obtained the results depicted in Fig. 14 and summarized in Table VII.
Figure 13: Track pattern adopted in the RT simulation to assess the channel characteristics between the original scenario and the simplified one.
Figure 14: NMSE CDF between the original scene and its simplified versions.
Environment Total duration (min) NMSE Number of faces
Average SD
Without simplification $94.89$ 5 922 92459229245\,922\,9245 922 924
Coverage Map cut-out $57.29$ $\mathbf{-13.49}$ $11.38$ 3 948 86539488653\,948\,8653 948 865
Sphere cut-out $42.18$ $-12.66$ $12.37$ 2 269 25022692502\,269\,2502 269 250
Rectangle cut-out $39.91$ $-10.70$ $0.251$ 947 22519472251\,947\,2251 947 225
Interactions cut-out $\mathbf{39.26}$ $-5.35$ $9.75$ 1 496 89214968921\,496\,8921 496 892
Vertex clustering $77.64$ $-3.80$ $8.20$ 1 812 24018122401\,812\,2401 812 240
Quadric edge collapse $77.62$ $-5.36$ $8.94$ 1 788 69217886921\,788\,6921 788 692
Figure 15: RT time simulation, in minutes, at Modern City scenario for each simplification method.
Furthermore, evaluating the RT simulation time on the same pre-defined path for each simplification method in comparison with the RT simulated in the original scenario. We obtained the results in Fig. 15 and Table VII, also considering the number of faces after the simplification. As the results show, among all simplification methods, those that presented the best results with respect to the trade-off between RT duration and NMSE are related to the Coverage Map and Sphere cut-out methods. The former obtained a better performance at maintaining channel characteristics, with a NMSE close to $-13.5$, as highlighted in Table VII. However, this method had a longer duration than other cut-out methods, such as the Sphere cut-out. The latter method obtained a slightly worse NMSE, equal to $-12.6$, but with a RT simulation time less than the former method. With respect to the other cut-out methods, we achieve the best RT duration with Interactions cut-out, but at the cost of an oversimplification and, consequently, an increase in the difference between the original scenario and the simplified one, since this method cuts out almost the entire scenario. Finally, the last cut-out method, the rectangle, had a similar RT duration to the interactions method, but with a lower NMSE, preserving the channel characteristics. Concerning the mesh simplification algorithms, i.e., the vertex clustering and the quadric edge collapse, both performed worse than the cut-out-based ones in terms of NMSE and RT duration. Both methods produced similar results, with the vertex clustering achieving a slightly better RT duration and the quadric edge collapse achieving a slightly better NMSE.
In the second experiment regarding Problem P2, the eight possible combinations of both simplification method approaches, along with the metric from the first experiment, produced the scatter plot that depicts solutions in a multi-objective Pareto front as shown in Fig. 16. This figure illustrates the tradeoff between RT simulation time (in minutes) on the $x$-axis and the average NMSE across all RT scenes on the $y$-axis. These results provide insight into the impact of combining different simplification methods, and are detailed in Table VIII. For instance, if a mesh simplification method is to be used, pairing it with the Coverage Map cut-out proves particularly effective, as it reduces simulation time while maintaining the NMSE. However, it is important to apply these combinations with caution, as they can lead to increased distortion.
@@
Figure 16: Tradeoff between NMSE and RT simulation time combining different solutions to Problem P2.
@@
TABLE VIII: Results obtained in terms of RT duration and NMSE for different combination of 3D simplification methods
@@
Rectangle cut-out + Edge collapse $42.19$ $-4.93$ $8.25$

### v2

@@ V-B2 CUT-OUT-BASED SIMPLIFICATION METHODS
  • Coverage Map Method: also requires an RT simulation and removes regions based on a predefined power threshold (in $\mathrm{d}\mathrm{B}$). The scenario is pruned by discarding areas where the received power falls below a threshold, optimizing the input for subsequent RT simulations.
@@ VI Evaluation
All experiments were performed using Sionna RT v0.19.20.19.20.19.20.19.2 [47] in different computers. For evaluating solutions to P1 (WCG methods), the server was equipped with an Intel® Core i7-10700F CPU @ 2.90 GHz, 128 GB RAM. For the P2 solutions (SOH methods), the server was equipped with Intel® Xeon Gold 5419Y, 32 GB RAM. In terms of channel models, we considered two types: a spatial 2D channel model and a 3D channel model [21, 22, 7]. The former model is defined by
$\mathbf{H}_{n}(f_{\textrm{s}})=\sum\limits_{r=1}^{K_{n}}\alpha_{r}e^{-j\Omega_{r}}\mathbf{a}_{\text{R}}(\theta_{r}^{\textrm{A}})\mathbf{a}_{\text{T}}^{*}(\theta_{r}^{\textrm{D}}),$ (5)
where $\mathbf{H}_{n}$ is the channel matrix for the scene $n$, $\mathbf{a}_{\text{T}}$, $\mathbf{a}_{\text{R}}$ are the steering vectors at the transmitter (Tx) and receiver (Rx); $\alpha_{r}$, $\theta_{r}^{\textrm{A}}$, $\theta_{r}^{\textrm{D}}$ and $\Omega_{r}$ are listed in Table I. Moreover, the phase $\Omega_{r}$ is the most difficult parameter to predict among the ones listed in Table I. It can be written as
$\Omega_{r}=\psi_{r}-2\pi f_{\textrm{s}}\tau_{r},$ (6)
where $\psi_{r}$ is the phase parcel that depends on the ray interactions and the electromagnetic properties (e. g., permittivity, conductivity, permeability, and scattering coefficient) of the respective materials [7], $f_{\textrm{s}}$ is the subcarrier frequency, and $\tau_{r}$ is the $r$-th path delay as listed in Table I.
Furthermore, the spatial 3D channel model is defined by
$\mathbf{H}_{n}(f_{\textrm{s}})=\sum\limits_{r=1}^{K_{n}}\alpha_{r}e^{-j\Omega_{r}}\mathbf{a}_{\text{R}}(\theta_{r}^{\textrm{A}},\phi_{r}^{\textrm{A}})\mathbf{a}_{\text{T}}^{*}(\theta_{r}^{\textrm{D}},\phi_{r}^{\textrm{D}}),$ (7)
where $\phi_{r}^{\textrm{A}}$ and $\phi_{r}^{\textrm{D}}$ are also listed in Table I.
Besides both WB channel models, in the evaluation of our proposed methods for P1, we also consider their NB variations, assuming a sufficiently small channel bandwidth $W$.
Finally, in order to evaluate the solutions to Problems P1 and P2, we consider each solution independently, in spite of the fact that in practice they can be combined by the WCG and SOH modules.
VI-A RT augmentation (P1)
To evaluate the solutions of problem P1, we considered the channel characteristics in both NB ($\tau_{r}W\ll 1$) and WB (which we adopted $W=100$ $\mathrm{M}\mathrm{H}\mathrm{z}$ and 64 subcarriers) regimes, based on Eq. (5) and Eq. (7). Specifically, two antenna array configurations were employed: (i) a MIMO system with a uniform linear array (ULA) comprising eight transmit and four receive antennas, and (ii) a MIMO system with a uniform rectangular array (URA) featuring a $4\times 4$ antenna layout in the Tx and $8\times 4$ in the Rx. The carrier frequency $f=2.14$ $\mathrm{G}\mathrm{H}\mathrm{z}$, an isotropic radiation pattern and cross-polarization were assumed in all configurations.
TABLE III: RT simulation parameters used to evaluate the proposed solution for the problem P1
Simulation parameter Parameter
@@ VI-A RT augmentation (P1)
Radiation pattern Isotropic
Polarization Cross polarization
@@
WB channel bandwidth ($\mathrm{M}\mathrm{H}\mathrm{z}$) 100
@@
$\hat{\mathbf{H}}_{p}=\mathbf{H}_{1}+\dfrac{p-1}{U}(\mathbf{H}_{2}-\mathbf{H}_{1}),$ (9)
where $1<p<U+1\hskip 4.26773pt\forall p\in\mathbb{Z}^{+}$.
To compare the performance of the ARTS method with this baseline in solving P1, we used three different urban macro scenarios: Munich, Etoile, and St. Canyon [47]. For all of them, the Tx height is greater than 25 $\mathrm{m}$. In terms of mobility, we have a single mobility condition, where only the Rx moves along a linear trajectory. For all Rx, we consider a constant velocity of 3 $\mathrm{k}\mathrm{m}\mathrm{/}\mathrm{h}$. The simulations consider only reflection and diffraction, and three different upsampling factors: $U=2$, $U=4$, and $U=10$. The value $U=2$ is the smallest upsampling factor, while $U=10$ is much more challenging. Given $U$, the sampling interval $T_{o}$, and the receiver speed, one can estimate the number of meters the receiver moved between neighboring scenes.
The conventional RT was executed for a total of $1000$ scenes, which are then processed and split into “known” channels and the ones that play the role of ground truth to assess augmentation. Table IV summarizes the experiments.
@@
# Array pattern $U$ Known scenes Known scenes density) $T_{\textrm{o}}$ $T_{\textrm{s}}$ Channel type
$1$ ULA $2$ $500$ $1$ scene every $0.4$ $\mathrm{m}$ $0.48$ $0.24$ NB
$2$ URA $2$ $500$ $1$ scene every $0.4$ $\mathrm{m}$ $0.48$ $0.24$ NB
$3$ ULA $10$ $100$ $1$ scene every $2$ $\mathrm{m}$ $2.4$ $0.24$ NB
$4$ ULA $4$ $250$ $1$ scene every $0.8$ $\mathrm{m}$ $0.96$ $0.24$ WB
TABLE V: Comparison of results with ARTS and the baseline for all four experiments
@@
Experiment 1 Matrix linear interp. $-1.40$ $5.56$ $-8.53$ $\mathbf{3.93}$ $-14.15$ $4.62$
Proposed ARTS method $\mathbf{-56.38}$ $\mathbf{4.44}$ $\mathbf{-67.26}$ $4.18$ $\mathbf{-75.91}$ $\mathbf{4.55}$
Experiment 2 Matrix linear interp. $1.72$ $5.10$ $-1.85$ $4.41$ $-2.95$ $9.15$
Proposed ARTS method $\mathbf{-55.39}$ $\mathbf{2.85}$ $\mathbf{-65.33}$ $\mathbf{4.36}$ $\mathbf{-70.10}$ $\mathbf{3.78}$
Experiment 3 Matrix linear interp. $-1.33$ $\mathbf{6.11}$ $-2.46$ $5.95$ $-5.16$ $11.31$
@@
Experiment 4 Matrix linear interp. $-0.43$ $\mathbf{5.94}$ $-5.74$ $\mathbf{4.62}$ $-9.81$ $\mathbf{7.73}$
Proposed ARTS method $\mathbf{-56.47}$ $9.34$ $\mathbf{-67.63}$ $5.22$ $\mathbf{-70.35}$ $14.79$
In the Experiments 1 and 2, $500$ channels are known, provided by the ray tracer, and augmentation is used to estimate the remaining $500$ NB channels, considering ULA and URA array patterns. The results are presented in Fig. 11, where the cumulative distribution functions shows that ARTS outperforms interpolation in the matrix domain.
Figure 11: CDF of NMSE for channel augmentation in ULA and URA MIMO channels, considering our proposed method for Problem P1 and matrix interpolation at Munich, Etoile, and St. Canyon scenario.
As mentioned, Experiment 3 is more challenging given that $U=10$. Specifically, it required generating $900$ scenes from information provided by $100$ known scenes. The results in Fig. 12 illustrate the performance under this configuration.
Figure 12: CDF of NMSE for channel augmentation in ULA MIMO channels considering our proposed method and matrix interpolation, with a fewer number of known samples.
The results in Fig. 12 indicate the expected decline in performance with respect to Fig. 11. This degradation can be attributed to the lack of spatial consistency between the RT samples. Specifically, the probability of obtaining two equivalent rays separated by 2 $\mathrm{m}$ is significantly lower compared to when the samples are only $0.4$ $\mathrm{m}$ apart. Despite this decline in NMSE, the proposed ARTS method outperforms the matrix interpolation method in all three scenarios, which is summarized in Table V. ARTS achieves a reduction of more than 17 $\mathrm{d}\mathrm{B}$ in average NMSE, while decreasing the maximum distortion in half of the cases.
Finally, the last set of experiments aimed to evaluate our method under a WB channel, which exhibits different characteristics from channel used in the previous three experiments, although conducted in the same scenarios. In this case, we employed an upsampling factor of $U=4$. The results, shown in Fig. 13, demonstrate that the ARTS method consistently outperforms the baseline under the WB channel model.
Figure 13: CDF of NMSE for channel augmentation in ULA WB MIMO channels, evaluated under the same scenarios as the previous three experiments.
@@
To validate the proposed solutions for the P2 problem, we consider only a ULA MIMO system, using the same number of Tx and Rx antennas as in P1, and adopt the NB scenario derived from Eq. 5. The carrier frequency, radiation pattern, and polarization remain unchanged. An important difference in the RT simulation, however, is that while the P1 solutions accounted only for reflection and diffraction, the P2 scenario also incorporates diffuse scattering. This inclusion increases the number of detected paths and, consequently, the complexity of the RT process, resulting in more pronounced effects on both NMSE and RT simulation time.
Using this channel configuration, we defined two experiments considering one more metric: the RT simulation time in comparison with that obtained in the original scenario. So, in the first experiment, we separately assessed all 3D simplification methods. In the second experiment, we evaluated the performance taking into account all possible combinations of mesh and cut-out-based simplification approaches. For both experiments, we considered the Modern City scenario, which was depicted in Fig. 2. Modern City has a larger number of meshes when compared to the three scenarios (Munich, Etoile, and St. Canyon) used to validate the WCG methods. More specifically, Modern City has 5 922 924 faces; while Munich, Etoile, and St. Canyon have 38 938, 13 058, and 74 faces, respectively. Furthermore, the scope of the two new experiments comprises six simplification methods, all defined in Fig. 8. Specific parameters for each method are summarized in Table VI.
@@
Environments Linear track Square track
RT duration (min) Average NMSE ($\mathrm{d}\mathrm{B}$) SD NMSE RT duration (min) Average NMSE ($\mathrm{d}\mathrm{B}$) SD NMSE
Without simplification $94.89$ $225.36$
Coverage Map cut-out $57.29$ $\mathbf{-13.48}$ $11.38$ $136.50$ $\mathbf{-7.80}$ $13.58$
Sphere cut-out $42.18$ $-12.40$ $11.38$ $76.10$ $-4.09$ $9.91$
Rectangle cut-out $39.91$ $-10.55$ $9.91$ $\mathbf{74.97}$ $-2.45$ $10.00$
Interactions cut-out $\mathbf{39.26}$ $-5.52$ $9.95$ $93.81$ $-5.11$ $10.96$
Vertex clustering $77.64$ $-3.76$ $7.96$ $188.80$ $-1.02$ $8.76$
Quadric edge collapse $77.62$ $-5.37$ $8.95$ $188.15$ $-2.23$ $10.60$
@@
To compute the NMSE and the RT runtime, for each simplification method, we considered two different track patterns: linear and square. For the linear track, we ran 162 scenes, varying the Rx. For the square track pattern, we considered a higher scene density, with a total of 284 scenes. Both tracks are shown in Fig. 14. For both cases, after that, we took the average, and the standard deviation (SD) between the NMSE values obtained. Thus, for this first experiment regarding P2, we obtained the results summarized in Table VII, divided by each track.
Figure 14: Two types of track mobility pattern adopted in the proposed solutions for P2 of the first experiment: linear (1) and square (2) track.
As the results show in Table VII, among all simplification methods, those that presented the best results with respect to the trade-off between RT duration and the average NMSE are related to the Coverage Map and Sphere cut-out methods. The former obtained a better performance at maintaining channel characteristics, with a NMSE close to $-13.5$ and $-7.8$, considering linear and square track mobility patterns, respectively. However, this method had a longer duration than other cut-out methods, such as the Sphere cut-out. The latter method obtained a slightly worse NMSE, equal to $-12.4$ in the linear track experiment, and $-4.09$ on the square track, but with a RT simulation time less than the former method. Among the different cut-out methods, the best RT runtime reduction was achieved using the Interactions cut-out on the linear track, with a time reduction of over 50%. However, this came at the cost of significant oversimplification, leading to a larger discrepancy between the channel characteristics of the original and simplified scenarios, as this method removes nearly the entire environment. A similar trend was observed with the Rectangle method on the square track, which also reduced RT duration but resulted in a less favorable NMSE. Conversely, on the linear track, the Rectangle method provided a substantial RT runtime reduction—close to that of the Interactions method—while offering a more balanced trade-off between complexity and accuracy.
Concerning the mesh simplification algorithms, i.e., the vertex clustering and the quadric edge collapse, both performed worse than the cut-out-based ones in terms of NMSE and RT duration. Both methods produced similar results, with the quadric edge collapse achieving a slightly better RT duration and average NMSE in both track patterns.
In the second experiment regarding Problem P2, using only the environment with a linear track, the eight possible combinations of both simplification method approaches, along with the metric from the first experiment, produced the scatter plot that depicts solutions in a multi-objective Pareto front as shown in Fig. 15.
@@
Figure 15: Tradeoff between NMSE and RT simulation time combining different P2 solutions considering a linear track pattern.
This figure illustrates the tradeoff between RT simulation time (in minutes) on the $x$-axis and the average NMSE across all RT scenes on the $y$-axis. These results provide insight into the impact of combining different simplification methods, and are detailed in Table VIII. For instance, if a mesh simplification method is to be used, pairing it with the Coverage Map cut-out proves particularly effective, as it reduces simulation time while maintaining the NMSE. However, it is important to apply these combinations with caution, as they can lead to increased distortion.
@@
TABLE VIII: Results obtained in terms of RT duration and NMSE for different combination of 3D simplification methods with Rx mobility assuming a linear track
@@
Rectangle cut-out + Edge collapse $42.19$ $\mathbf{-4.93}$ $8.25$